sin (tan -1 x)/1 + x 2 please solve the above question

sin (tan-1 x)/1 + x2


please solve the above question


 

Grade:12

6 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

dear Shinoy

put tan-1x = t

(1/1+x2)dx =dt

put this

integral becomes ∫sint dt = -cost + C

                                  = -cos(tan-1x) + C   Answer

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

aditya suman
13 years ago

put tan-1x=t

then , dt/dx=1/1+x2

dt=dx/1+x2                _1

now ,sintdx/1+xusing 1

=sintdt=-cost +c=-cos(tan-1x)+c

vikas askiitian expert
13 years ago

 I=(sintan-1x) /1+x2

put tan-1x =t

   differentiating wrt x

1/1+x2 dx =dt

I=sintdt=-cost + c

  =-costan-1x + c

arvind ramesh
13 years ago

sub tan-1x as t....dt is 1/1+x^2 dx =>intg sint dt =>ans.= -cos(tan-1x)+c;

rajan jha
13 years ago

let (tan-1 x)=t

 

then dt/dxsqrt()=1/1 + x2

then the integration becomes

 sin(t)dt =-cos(t)

=-cos((tan-1 x))

=-cos(sec-1sqrt(1 + x2))

= -cos(cos-1sqrt(1/1+x2))

=-sqrt((1/1+x2).

hence the answer

like my answer?

Souradeep Majumder
13 years ago

put tan-1x =z..  then sin-1x dx =dz.  thaen it is in the form integration sinzdz and ans is -cosz+c put z=tan-1x and you will find the value  -cos(tan-1x)+c.

 

approve my answer..................